OpenStudy (elisaneedshelp):

Multiplying rational expressions involving multivariate monomials, I'll give you a medal Please

OpenStudy (elisaneedshelp):

\[ \frac{ 6x}{ 2b}\times \frac{ 8bx }{ 3x}\]

OpenStudy (elisaneedshelp):

I understand the basic concept of what to do but I keep getting it wrong since every problem has its little difference

OpenStudy (jhannybean):

How did you initially go about solving it?

OpenStudy (elisaneedshelp):

well I multiply and I get 48/6 and i cancel out the variables this time?

OpenStudy (jhannybean):

Well, if you treat \(x\), and \(b\), both as variables, would you not be able to cancel them out via cross multiplication?

OpenStudy (jhannybean):

Cancel like terms with like terms

OpenStudy (elisaneedshelp):

Can we try another one ? I would continue with this one but my electronic workbook kicked me off that problem

OpenStudy (jhannybean):

\[\huge \frac{\color{orange}{6}\color{red}{x}}{\color{orange}{2}\color{blue}{b}} \cdot \frac{\color{orange}{8}\color{blue}{b}\color{red}{x}}{\color{orange}{3}\color{red}{x}}\]

OpenStudy (jhannybean):

Sure.

OpenStudy (elisaneedshelp):

Sorry and thanks.\[\frac{ 3a }{ 2x}\times \frac{ 4x ^{5} }{ 9ax}\]

OpenStudy (elisaneedshelp):

That is a 4x^5

OpenStudy (jhannybean):

\[\huge \frac{\color{orange}{3}\color{blue}{a}}{\color{orange}{2}\color{red}{x}} \cdot \frac{\color{orange}{4}\color{red}{x^5}}{\color{orange}{9}\color{blue}{a}\color{red}{x}}\]

OpenStudy (jhannybean):

so comparing all the like colored terms together, see how we can cross cancel the blue a terms?

OpenStudy (elisaneedshelp):

is it okay if I multiply first and then cross them out? or I shouldnt do that in the future?

OpenStudy (elisaneedshelp):

so far I have \[\frac{ 12 }{ 18x ^{3} }\]

OpenStudy (jhannybean):

it's easier to cross cancel terms that can be simplified first because sometimes you'll get gigantic numbers that will overwhelm you.

OpenStudy (jhannybean):

if you're multiplying them all out, what happened to your \(a\) terms?

OpenStudy (jhannybean):

Oh I see, so you've partially reduced your function.

OpenStudy (jhannybean):

So how I would solve it is just cross cancel everything I can simplify: \[\huge \frac{\cancel{\color{orange}{3}}\color{blue}{a}}{\cancel{\color{orange}{2}}\color{red}{x}} \cdot \frac{2\cancel{\color{orange}{4}}\color{red}{x^5}}{3\cancel{\color{orange}{9}}\color{blue}{a}\color{red}{x}}\] this leaves me with: \[\huge \frac{2\color{blue}{a}\color{red}{x^5}}{3\color{blue}{a}\color{red}{x^2}}\]Now i can cancel out by a's and i'll be left with : \[\huge \frac{2\color{red}{x^5}}{3\color{red}{x^2}}\]using the power subtraction rule for exponents, I will end up with my final answer being :\[\huge \frac{2}{3}x^{5-2} = \frac{2}{3} x^3\]

OpenStudy (elisaneedshelp):

Wow and could the same thing be done for \[\frac{ 15y }{ 2b ^{2} }\times \frac{ b ^{3} y}{ 9y ^{4} }\]

OpenStudy (elisaneedshelp):

well not exactly the same which is what i hate

OpenStudy (jhannybean):

It will work the same way :) yes!

OpenStudy (elisaneedshelp):

let me try now

OpenStudy (elisaneedshelp):

do I simplify before I multiply?

OpenStudy (jhannybean):

Yes :)

OpenStudy (elisaneedshelp):

Do I only simplify the 15 and the 9 into 3s? or am i completely wrong?

OpenStudy (elisaneedshelp):

no the 15 into a 5 and the 9 into a 3

OpenStudy (elisaneedshelp):

I need guidance I'm confused especially with the variables

OpenStudy (jhannybean):

yes you're correct, simplify them both by the LCM

OpenStudy (jhannybean):

\[\huge \frac{ \color{red}{5}\cancel{15}y }{ 2b ^{2} }\times \frac{ b ^{3} y}{ \color{red}{3}\cancel{9}y ^{4} }\]

OpenStudy (elisaneedshelp):

I also have confusion on where to place each variable on the final answer. I have plain b and y^6

OpenStudy (jhannybean):

You keep the variables where they are unless they can be reduced and canceled out, similar to our last problem if you scrolled up ^^

OpenStudy (jhannybean):

So with this problem we can do one of two things: 1. cross multiply and cancel out the \(y\)'s that can be cancelled 2. multiply across the numerator and denominator and worry about cancelling afterward Which would you prefer?

OpenStudy (elisaneedshelp):

Lol but the variables are everywhere. Maybe number 2 is the best for me

OpenStudy (jhannybean):

Alright

OpenStudy (jhannybean):

Now that we've cross multiplied and reduced the numbers, let's multiply it out :) \[\huge \frac{5y^2b^3}{6b^2y^4}\] I didnt see the other b there at first! :o Do you see how I got this fraction? ^^

OpenStudy (elisaneedshelp):

Yeah I see Thats my safe zone

OpenStudy (jhannybean):

:) alright

OpenStudy (jhannybean):

Now let's separate each like term into a fraction, this will help us use the subtraction rule for exponents \[\huge \frac{5}{6} \cdot \frac{b^3}{b^2} \cdot \frac{y^2}{y^4}\] Are you comfortable seeign it like this? :o

OpenStudy (elisaneedshelp):

yes :)

OpenStudy (jhannybean):

Now we apply the subtraction rule for exponents, and we leave whatever fractions that cannot be simplified any further alone. \[\huge \frac{5}{6} \cdot b^{3-2} \cdot y^{2-4}\] What does this give you when simplified, can you tell me? :P

OpenStudy (elisaneedshelp):

b and y^-2

OpenStudy (jhannybean):

Good. We want t write al our powers as positive powers, therefore we have to change \(y^{-2}\) to \(\dfrac{1}{y^2}\) to keep it positive. :) Following?

OpenStudy (elisaneedshelp):

Yeap following

OpenStudy (jhannybean):

therefore we can rewrite our function as: \[\huge \frac{5}{6} \cdot b \cdot \frac{1}{y^2} = \frac{5b}{6y^2}\]

OpenStudy (elisaneedshelp):

I got your message but OS wont let me reply but its alright :P lol

OpenStudy (jhannybean):

Do you understand this better, now? :D

OpenStudy (elisaneedshelp):

Yes I'll try doing the other one completely by myself!

OpenStudy (jhannybean):

Yay!

OpenStudy (elisaneedshelp):

Am I right so far? on #3

OpenStudy (elisaneedshelp):

Lol i tried to do it all by myself

OpenStudy (elisaneedshelp):

is the final answer\[\frac{ 12b ^{2} }{ y ^{4} }\]

OpenStudy (elisaneedshelp):

No its not but i tried

OpenStudy (jhannybean):

number 2 seems wrong.... we worked this out just now though ^^

OpenStudy (elisaneedshelp):

Lol yeah dont worry I havent erased but I will go through our thread and rewrite it

OpenStudy (jhannybean):

Oh okay

OpenStudy (jhannybean):

\[\begin{align} \huge \frac{8b^3y^4}{3y} \cdot \frac{9y}{2b} &= \huge \frac{72}{6}\cdot \frac{b^3}{b}\cdot \frac{y^5}{y} \\ &= \huge \frac{72}{6} \cdot b^{3-1} \cdot y^{5-1} \\&= \huge ....?\end{align}\]

OpenStudy (jhannybean):

Try working it out this way, and see if this method is easier for you. I've got to head off OS right now, so good luck on solving the rest!!

OpenStudy (elisaneedshelp):

Thanks for the help Jhannybean ! :)

OpenStudy (jhannybean):

no problem:)